# mystyle backgroundcolor=white!10

### v1

@@ III-A Sufficient Conditions for Upper Bounds
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (3)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\gamma\in[0,1]\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha\in(0,1)\wedge\gamma\in[0,1]\\ 1-(1-v(\bm{x}_{0}))(1-\beta)^{N}&\text{if~{}}\alpha\in(0,1]\wedge\gamma\in(-\infty,0)\end{cases},$
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\widetilde{\mathcal{X}}\\ v(\bm{x})\geq 1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})&\forall\bm{x}\in\widetilde{\mathcal{X}}\end{cases}.$ (4)
@@
3) We will show the case of $\alpha\in(0,1]\wedge\gamma\in(\infty,0)$.
Since $\beta<1-\frac{1}{\alpha}$, thus $\frac{1}{\alpha}<1-\beta$. Consequently, $01-\beta>$ 0. Thus, we conclude that if $v(\bm{x})$ satisfies (3), it will satisfy
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq v(\bm{x})(1-\beta)+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\leq(1-\beta)v(\bm{x})+\beta&\forall\bm{x}\in\widetilde{\mathcal{X}}\\ v(\bm{x})\geq 1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})&\forall\bm{x}\in\widetilde{\mathcal{X}}\end{cases}.$
@@
If $\alpha=1$ and $\beta=0$, Theorem 1 is equivalent to Proposition 3 in [16], which formulates a sufficient condition for determining an upper bound of the probability of exiting the safe set $\mathcal{X}$ eventually(i.e., $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}$). The upper bound is $v(\bm{x}_{0})$, which is also an upper bound of the exit probability $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}$. Besides, since $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=\alpha^{-N}(v(\bm{x}_{0})-\frac{\alpha\beta}{\alpha-1})+\frac{\alpha\beta}{\alpha-1}$, we conclude that when $\gamma\in[0,1]$ and $\alpha\in(0,1)$, if $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$, $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$ is also an upper bound of the probability of exiting the safe set $\mathcal{X}$ eventually; if $v(\bm{x}_{0})<\frac{\alpha\beta}{\alpha-1}$,
$\lim_{N\rightarrow\infty}v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=-\infty,$
which implies $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$ and thus $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$. Similarly, when $\alpha\in(0,1)$ and $\gamma\in(-\infty,0)$, if $v(\bm{x}_{0})<1$, $\lim_{N\rightarrow\infty}1-(1-v(\bm{x}_{0}))(1-\beta)^{N}=-\infty$ holds, implying $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$ and $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$.
Theorem 1 extends the conclusion drawn in [6], as it allows for a value of $\alpha$ within the interval $(0,1)$, in contrast to the requirement of $\alpha\geq 1$ in the cited work. For convenience of reference, we present the related result in [6] here. If there exists a continuous nonnegative function satisfying
$\begin{cases}v(\bm{x}_{0})<1&\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (5)
$\alpha\geq 1$, and $\beta\geq 0$, then $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\leq$
$\begin{cases}1-(1-v(\bm{x}_{0}))(1-\beta)^{N}&\text{if~{}}\alpha>1\wedge\gamma\in(-\infty,0]\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha>1\wedge\gamma\in(0,\infty)\\ v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\gamma\in(0,\infty)\end{cases},$
@@
When comparing constraints (3) and (5), it becomes apparent that if there exists a function $v(\bm{x})$ along with $\alpha>1$ and $\beta\in[0,1]$ satisfying (5), the same function $v(\bm{x})$ with $\frac{1}{\alpha}$ and $\beta\in[0,1]$ will satisfy (3). Consequently, is can be concluded that (3) is more expressive than (5).
III-B Sufficient Conditions for Lower Bounds
@@
Theorem 2
@@
$\begin{cases}\beta+\alpha v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (6)
@@
$\begin{cases}\frac{(\alpha^{N+1}v(\bm{x}_{0})-M)(\alpha-1)+\beta(\alpha^{N+1}-1)}{(\alpha+\beta-1)(\alpha^{N+1}-1)}&\text{if~{}}\alpha>1\\ 1+\frac{v(\bm{x}_{0})-M}{\beta(N+1)}&\text{if~{}}\alpha=1\end{cases}.$
Proof:
@@
The last inequality is obtained according to Lemma 1, which states $\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\geq\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}$ for $i\leq N$, and the fact that $01-\alpha-\beta<$ 0.
@@
It is worth noting here that if there exists a bounded function $v(\bm{x})\colon\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying (6) with $\alpha=1$, then system (1), starting from $\bm{x}_{0}\in\mathcal{X}$, will exit the safe set $\mathcal{X}$ eventually with the probability of one, since $\lim_{N\rightarrow\mathbb{N}}1+\frac{v(\bm{x}_{0})-M}{\beta(N+1)}=1$. Therefore, if system (1) does not feature this property, $\alpha=1$ cannot be used to perform computations.
Remark 1
@@
Remark 2
One may wander whether a sufficient condition for lower bounding the probability of exiting the safe set $\mathcal{X}$ over the time horizon $[0,N]$ in Problem 1 can be constructed by directly reversing the sign in constraint (3) in Theorem 1. We will give a brief explanation here that the condition constructed in this manner will consistently yield zero as lower bounds.
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\geq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (7)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\gamma\in(-\infty,0)\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha\in(0,1)\wedge\gamma\in(-\infty,0)\\ 1-(1-v(\bm{x}_{0}))\alpha^{-N}&\text{if~{}}\alpha\in(0,1]\wedge\gamma\in[0,\infty)\end{cases},$
@@
IV Finite-time Reach-avoid Verification
In this section, we present our sufficient conditions for characterizing upper and lower bounds of the probability in Problem 2. The sufficient condition for upper bounds is formulated in Subsection IV-A and the one for lower bounds is introduced in Subsection IV-B.

### v2

@@ 3.1 Sufficient Conditions for Upper Bounds
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta,&\forall\bm{x}\in\mathcal{X},\\ v(\bm{x})\geq 1,&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X},\\ v(\bm{x})\geq 0,&\forall\bm{x}\in\mathcal{X},\end{cases}$ (3)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N,&\text{if~}\alpha=1\wedge\gamma\in[0,1],\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1},&\text{if~}\alpha\in(0,1)\wedge\gamma\in[0,1],\\ 1-(1-v(\bm{x}_{0}))(1-\beta)^{N},&\text{if~}\alpha\in(0,1]\wedge\gamma\in(-\infty,0),\end{cases}$
@@
[…]
@@
3) We will show the case of $\alpha\in(0,1]\wedge\gamma\in(-\infty,0)$.
Since $\beta<1-\frac{1}{\alpha}$, thus $\frac{1}{\alpha}<1-\beta$. Consequently, $1-\beta>0$. Thus, we conclude that if $v(\bm{x})$ satisfies (3), it will satisfy
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq v(\bm{x})(1-\beta)+\beta,&\forall\bm{x}\in\mathcal{X},\\ v(\bm{x})\geq 1,&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X},\\ v(\bm{x})\geq 0,&\forall\bm{x}\in\mathcal{X},\end{cases}$
@@
[…]
@@
If $\alpha=1$ and $\beta=0$, Theorem 1 is equivalent to Proposition 3 in [37], which formulates a sufficient condition for determining an upper bound of the probability of exiting the safe set $\mathcal{X}$ eventually (i.e., $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)$). The upper bound is $v(\bm{x}_{0})$, which is also an upper bound of the exit probability $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)$. On the other hand, when $\alpha\in(0,1)$ and $\gamma\in(-\infty,0)$, if $v(\bm{x}_{0})<1$, $\lim_{N\rightarrow+\infty}1-(1-v(\bm{x}_{0}))(1-\beta)^{N}=$-$\infty$ holds, implying $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)=0$ and $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)=0$.
Theorem 1 extends the conclusion drawn in [17], as it allows for a value of $\alpha$ within the interval $(0,1)$, in contrast to the requirement of $\alpha\geq 1$ in the cited work. For convenience of reference, we present the related result in [17] here, which corresponds to Theorem 3 in Chapter 3 in [17].
Proposition 1.
If there exists a continuous nonnegative function $v(\cdot):\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying
$\begin{cases}v(\bm{x}_{0})<1,&\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta,&\forall\bm{x}\in\mathcal{X},\\ v(\bm{x})\geq 1,&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X},\\ v(\bm{x})\geq 0,&\forall\bm{x}\in\mathcal{X},\end{cases}$ (5)
where $\alpha\geq 1$ and $\beta\geq 0$, then $\mathbb{P}^{\infty}\big(\forall k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)\leq$
$\begin{cases}1-(1-v(\bm{x}_{0}))(1-\beta)^{N},&\text{if~}\alpha>1\wedge\gamma\in(-\infty,0],\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1},&\text{if~}\alpha>1\wedge\gamma\in(0,\infty),\\ v(\bm{x}_{0})+\beta N,&\text{if~}\alpha=1\wedge\gamma\in(0,\infty),\end{cases}$
@@
When comparing constraints (3) and (5), it becomes apparent that if there exists a function $v(\bm{x})$ along with $\alpha>1$ and $\beta\in[0,1]$ satisfying (5), the same function $v(\bm{x})$ with $\frac{1}{\alpha}$ and $\beta\in[0,1]$ will satisfy (3). In addition, it is observed that the barrier function in [20], referred to as c-martingales, is just a function that satisfies constraint (3) with $\alpha=1$ and $\beta\geq 0$.
3.2 Sufficient Conditions for Lower Bounds
@@
The sufficient condition requires the existence of a function $v(\bm{x})$ defined on the set $\widetilde{\mathcal{X}}$ that satisfies three key properties: (1) it admits a finite upper bound over $\widetilde{\mathcal{X}}$, (2) it is less than or equal to one on the subset $\widetilde{\mathcal{X}}\setminus\mathcal{X}$, and (3) its expected value at the next step along the system dynamics described in (1) increases over its $\alpha$-scaled current value by more than $\beta$ for $\bm{x}\in\mathcal{X}$, where $\alpha\in[1,\infty)$, and $\beta\in(1-\alpha,\infty)$. The lower bounds derivation proceeds as follows: Lemma 1 establishes that the probability of system (2) entering $\widetilde{\mathcal{X}}\setminus\mathcal{X}$ at time $t=N$ is equal to the probability of system (1) entering this set up to time $t=N$. Then, we derive the lower bounds using system (2) by analyzing $\alpha$. Relying on the fact that the system in (2) remains stationary when starting from $\widetilde{\mathcal{X}}\setminus\mathcal{X}$, which implies $\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}$, and noting that $v(\bm{x})$ is less than or equal to one on this subset, we reformulate the constraint on $v(\bm{x})$. Specifically, using system (2), we require that the expected value of $v(\bm{x})$ at the next step grows at least by $\beta-(\alpha-1+\beta)1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})$ over its $\alpha$-scaled current value $v(\bm{x})$, for all $\bm{x}\in\widetilde{\mathcal{X}}$. For $\alpha>1$ and $\alpha=1$, the lower bounds are subsequently obtained through recursive application of the reformulated constraint, in conjunction with the upper bound on the function $v(\bm{x}):\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$.
Theorem 2.
@@
$\begin{cases}\beta+\alpha v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))],&\forall\bm{x}\in\mathcal{X},\\ v(\bm{x})\leq 1,&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X},\end{cases}$ (6)
@@
$\begin{cases}\frac{(\alpha^{N+1}v(\bm{x}_{0})-M)(\alpha-1)+\beta(\alpha^{N+1}-1)}{(\alpha+\beta-1)(\alpha^{N+1}-1)},&\text{if~}\alpha>1,\\ 1+\frac{v(\bm{x}_{0})-M}{\beta(N+1)},&\text{if~}\alpha=1.\end{cases}$
Proof.
1) We first prove the case of $\alpha>1$.
@@
The last inequality is obtained according to Lemma 1, which states $\mathbb{P}^{\infty}\big(\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)\geq\mathbb{P}^{\infty}\big(\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big)$ for $i\leq N$, and the fact that $1-\alpha-\beta<0$.
@@
2) The conclusion for the case of $\alpha=1$ can be justified via following the proof of the above one.
@@
It is worth noting here that if there exists a bounded function $v(\bm{x})\colon\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying (6) with $\alpha=1$, then system (1), starting from $\bm{x}_{0}\in\mathcal{X}$, will exit the safe set $\mathcal{X}$ eventually with the probability of one, since $\lim_{N\rightarrow+\infty}1+\frac{v(\bm{x}_{0})-M}{\beta(N+1)}=1$. Therefore, if system (1) does not feature this property, $\alpha=1$ cannot be used to perform computations.
Remark 1.
@@
Remark 2.
One may wonder whether a sufficient condition for lower bounding the probability of exiting the safe set $\mathcal{X}$ over the time horizon $[0,N]$ in Problem 1 can be constructed by directly reversing the sign in constraint (3) in Theorem 1. We will give a brief explanation here that the condition constructed in this manner will consistently yield zero as lower bounds.
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\geq\frac{v(\bm{x})}{\alpha}+\beta,&\forall\bm{x}\in\mathcal{X},\\ v(\bm{x})\leq 1,&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X},\\ v(\bm{x})\leq 0,&\forall\bm{x}\in\mathcal{X},\end{cases}$ (7)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N,&\text{if~}\alpha=1\wedge\gamma\in(-\infty,0),\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1},&\text{if~}\alpha\in(0,1)\wedge\gamma\in(-\infty,0),\\ 1-(1-v(\bm{x}_{0}))\alpha^{-N},&\text{if~}\alpha\in(0,1]\wedge\gamma\in[0,\infty),\end{cases}$
@@
4 Finite-time Reach-avoid Verification
In this section, we present our sufficient conditions for characterizing upper and lower bounds of the reach-avoid probability in Problem 2. The sufficient condition for upper bounds is formulated in Subsection 4.1 and the one for lower bounds is introduced in Subsection 4.2.